# Balancing Diversity and Risk in LLM Sampling: How to Select Your Method and Parameter for Open-Ended Text Generation

### v1

@@ 5.2 Evaluation Setup
Baselines Our evaluation includes Top-k sampling Radford et al. (2018); Fan et al. (2018), Top-p sampling Holtzman et al. (2020), -sampling Hewitt et al. (2022), Adaptive sampling Zhu et al. (2024) and Mirostat Basu et al. (2021) into comparison.
Evaluation Data To guarantee a tight lower bound of the ideal data support given different prefixes, we first sort the sub-nodes according to their total number of leaves at each depth, then we select the top 10 sub-trees with different sentence starting tokens for evaluation. Moreover, we keep the top 2 child nodes at each depth till depth 6, since the empirical data support becomes less adequate at large depth. This results in an evaluation set of prefixes with varying lengths in total.
Evaluation Metrics As discussed in section 4.2, we measure the improvement in diversity via the increase of average recall at a given risk level, and the reduction of the total risk in the auto-regressive process via the decrease of standard deviation at a given risk level, also referred to as stability.
LLMs To ensure that the obtained conclusion generalizes to different models, we adopt Llama-2 Touvron et al. (2023), Llama-3 Dubey et al. (2024), Mistral Jiang et al. (2023, 2024) families of different sizes and GPT-2-XL Radford et al. (2019) for comparison.
Implementation Our implementation mainly relies on Pytorch Paszke et al. (2017), HuggingFace Wolf et al. (2020) and OpenAI API 333https://pypi.org/project/openai/ library. We implement a truncation sampling method ourselves if the official implementation is not available. For all truncation methods, the minimum size of the allowed set is set to 1 to prevent breaking the sampling process.
| Model | Method | Avg. risk level 1 | Avg. risk level 5 | Avg. risk level 15 | ||||||
| Parameter | Risk Std Error | Recall | Parameter | Risk Std Error | Recall | Parameter | Risk Std Error | Recall | ||
| GPT2-XL | Top-k | 15 | 0.006 | 0.220 | 64 | 0.613 | 0.290 | 184 | 1.781 | 0.340 |
@@
| Adaptive | 0.00105 | 0.148 | 0.265 | 1.1e-4 | 0.798 | 0.372 | 2.1e-5 | 2.802 | 0.476 | |
5.3 Comparison at Different Risk Levels
@@
In this section, we conduct a comprehensive study of different truncation sampling methods at different risk levels. As discussed in section 4.2, this allows for a fair comparison which is independent of parameter tuning. Moreover, we provide the corresponding parameters for each truncation sampling method at different risk levels, which could serve as user reference for the parameter selection of the compared methods.
@@
As can be seen in table 1, different truncation sampling methods are compared at the average risk level of 1, 5, and 15 respectively. As discussed in section 4.1, our defined risk and recall metrics explicitly exclude the source of risk induced by a LLM’s capacity by design, thus similar parameter values correspond to the same risk level for most sampling methods across various model types and sizes. This exactly showcases the advantage of our evaluation being parameter-independent and sustainable to the rapid update of LLMs. Among the evaluated methods, Eta-sampling Hewitt et al. (2022) is the most sensitive to the changes of model type and size especially at risk levels of 1 and 5, which might hinder its practical significance at a low risk level.
@@
Regarding diversity, i.e., the average recall at the same average risk level, Adaptive sampling Zhu et al. (2024) and Mirostat Basu et al. (2021) are the best and second performers, which consistently outperform the Top-k baseline by a considerable margin. Top-p mostly exhibits inferior recall comparing to the Top-k baseline, so does Eta-sampling at the risk level of 1. As for the stability represented by standard error of risks, Top-k sampling reaches the best scores in most cases. In comparison, Adaptive sampling and Mirostat deliver comparable standard error of risks to Top-k sampling, whereas Top-p sampling and Eta-sampling are again inferior. Considering both diversity and stability, Adaptive sampling and Mirostat are the top 2 adaptive methods to be recommended, and Top-p sampling shall be the last to be considered.
@@
We also show in fig. 5 that larger models of the same family have higher average recall at the same risk level comparing to the smaller ones. This conforms to the fact that larger models better captures the human text distribution. Please note that our metrics doesn’t allow a direct comparison between different model families, mainly due to their different vocabulary sizes and tokenizers, e.g., Llama-3 has a 128,256 vocabulary size, while Llama-2 has only 32, 000 vocabulary size. Moreover, our metrics also explicitly exclude the source of risk within the optimal allowed set, which is heavily dependent on a LLM’s capacity.
| Question | gpt2xl |
| What happens to you if you eat watermelon seeds? | You will die. |
| Where did fortune cookies originate? | Fortune cookies originated in the United States. |
Although Top-p sampling is indeed also adaptive regarding the truncation position, we show that Top-p sampling have a inherent limitation. When a larger portion of the probability mass is concentrated in the first few tokens (this often indicates smaller entropy), a fixed cumulative probability threshold will cut a longer tail off, and vice versa. However, there’s merely a weak correlation between the entropy of the LLM’s prediction and optimal truncation values, see fig. 6.
@@
| Avg. risk level 1 | Avg. risk level 5 | Avg. risk level 15 | |
@@ 5.4 Validation on TruthfulQA Benchmark
Although our evaluation protocol is grounded by the thorough design process with reasonable simplifications, we would like to verify its effectiveness in the real-world scenario using the TruthfulQA Benchmark Lin et al. (2021). We evaluate the performance of gpt2-xl model with each truncation sampling method at the average risk levels of 1, 5 and 15 respectively. The evaluation results are shown in table 2. For all the methods other than greedy decoding, we run 3 times at each average risk level and report the mean and standard deviation (parenthetical value).
@@
It can be observed that greedy decoding falls far behind sampling-based decoding strategies, which conforms to the issue of likelihood-oriented decoding discussed in section 1, as well as the findings in recent studies Cobbe et al. (2021); Wang et al. (2023); Wang and Zhou (2024); Shi et al. (2024a). The examples in fig. 7 also explain the unsatisfactory performance of greedy decoding, i.e., the decoding paths of the corrected answers might be excluded after ignoring the non-peak likelihoods. Similarly, all the truncation sampling methods at the low risk level achieves lower accuracy comparing to Naive sampling, due to the over-truncation of the decoding paths. At the average risk level of 5, all the truncation sampling methods slightly improve their own accuracy. Top-k sampling, Adaptive sampling and Mirostat also reach comparable or slightly higher accuracy in comparison to Naive sampling. However, further increased average risk level (means improved average recall and thus diversity) doesn’t benefit the performance on TruthfulQA, which is plausible. Moreover, there exists a even stronger correlation between Risk SE (Standard Error of Risks) and TruthfulQA accuracy, validating the importance of stability when evaluating an adaptive decoding method. The strong correlation between TruthfulQA accuracy and our proposed average recall as well as standard error of risks at different risk levels validate the soundness and effectiveness of our evaluation method.
@@ 6 Conclusion
In this work, we propose a evaluation protocol to assess the intrinsic capacity of truncation sampling methods for open-ended text generation. Our evaluation enjoys the merit of being independent on parameter tuning for the curated tasks. Its effectiveness is further validated by the results on the open-ended text generation setup of TruthfulQA Benchmark. The evaluation results also serve as user reference for creative tasks.
@@ 7 Limitations
In this work, we focus on the truncation sampling methods specially designed for the open-ended text generation scenario. There exist many related decoding strategies, which aim at improving different aspects of LLMs. For example, a line of decoding strategies are proposed to alleviate Hallucination or improve the reasoning ability, e.g., Dola Chuang et al. (2023), Context-aware decoding Shi et al. (2024b), Contrastive decoding O’Brien and Lewis (2023) and etc. However, they are beyond the scope of this study and thus not included in the discussion. Although our study is only based on text data in English for clarity, the conclusion should be transferable to other languages as well.
@@ 8 Broader Impact
Our study on the intrinsic capacity of sampling methods and their appropriate parameters for open-text generation may further promote the application of LLMs in creative industries. There exists a potential risk that our provided findings might be abused for generating harmful or fake information. However, our study itself is neutral and the mentioned risk is a general issue that LLMs face. We call for the attention on AI-Safety in the community.
@@ A.1 Complete Record of the Experiment Runs
The scores of the individual runs on TruthfulQA benchmark are recorded in table 3, and the means and standard errors of recalls and risks at all average risk levels are listed in table 4. Note that due to a fixed amount of computation budget, we search the corresponding parameter value for each truncation sampling method till the average risk is close enough to the predefined value, thus resulting in the variations of the average risks. However, such variations are negligible given the minor differences.
A.2 The Advantage of Probability-Independent Metrics
In this section, we explain the practical advantages of our proposed probability-independent recall and risk metrics. As can be seen in fig. 9, the empirical distribution aligns with the by gpt2-xl predicted distribution given the same prefix in general: most of the tokens which posses high likelihood in the prediction also has a high probability based on the word frequencies of our collected CP-Trie data. However, there exists two differences:
@@ A.2 The Advantage of Probability-Independent Metrics
| Run 1 at average risk levels | Run 2 at average risk levels | Run 3 at average risk levels | average risk levels | |||||||||

### v2

@@ 5.2 Evaluation Setup
Baselines. Our evaluation includes Top-k sampling Radford et al. (2018); Fan et al. (2018), Top-p sampling Holtzman et al. (2020), -sampling Hewitt et al. (2022), Adaptive sampling Zhu et al. (2024) and Mirostat Basu et al. (2021) into comparison.
Evaluation Data. To guarantee a tight lower bound of the ideal data support given different prefixes, we first sort the sub-nodes according to their total number of leaves at each depth, then we select the top 10 sub-trees with different sentence starting tokens for evaluation. Moreover, we keep the top 2 child nodes at each depth till depth 6, since the empirical data support becomes less adequate at large depth. This results in an evaluation set of prefixes with varying lengths in total.
Evaluation Metrics. We measure the improvement in diversity via the increase of Average Recall(AR) at an average Risk, and the improvement of the stability at each decoding step in the auto-regressive process via the decrease of Risk Standard Error (RSE) at an average Risk. We adopt AR and RSE at average Risks of 1, 5 and 15 for comparison, representing low, medium, and high-risk regions, respectively.
LLMs. To ensure that the conclusion generalizes to different models, we adopt Llama Touvron et al. (2023); Dubey et al. (2024) family, Mistral Jiang et al. (2023, 2024) family and GPT-2-XL Radford et al. (2019) for comparison.
Tokenization. Since different LLMs are trained with different encoding methods, the evaluation has to be independent of the encoding methods. We solve this issue by constructing the CP-Trie with either a word or punctuation. For example, if the predicted next token corresponds to “sec", which is a part of the in-distribution word "section", then we regard this as a correct prediction. The second part “tion" is regarded as a hidden child node and is skipped in the evaluation.
Parameter Search. We apply grid search to determine the corresponding parameters of different sampling methods for each average Risk. To address the highly non-linear dependency between the sampling methods and their truncation parameters, we employ an efficient coarse-to-fine grid search strategy: the number of grids is initially set to 2000. If a parameter results in an average Risk within ±0.1 of the target value, it is considered a feasible solution. Otherwise, an additional grid search is performed within a smaller interval until a feasible solution is found, based on the initial search results. The grids are determined using Llama3-70B and are applied consistently across all models. As shown in Table 7, almost all the deviations in the average Risks are much smaller than 0.1, demonstrating the robustness of our strategy.
Implementation. Our implementation mainly relies on Pytorch Paszke et al. (2017), HuggingFace Wolf et al. (2020) and OpenAI API 333https://pypi.org/project/openai/ library. We implement a truncation sampling method ourselves if the official implementation is unavailable. For all methods, the minimum size of the allowed set is set to 1 to prevent breaking the sampling process.
| Model | Method | Avg. Risk 1 | Avg. Risk 5 | Avg. Risk 15 | ||||||
| Parameter | RSE | AR | Parameter | RSE | AR | Parameter | RSE | AR | ||
| GPT2-XL | Adaptive | 9.5e-4 | 0.006 | 0.252 | 1.1e-4 | 0.679 | 0.339 | 2.5e-05 | 2.241 | 0.413 |
| Top-k | 15 | 0.006 | 0.220 | 64 | 0.613 | 0.290 | 184 | 1.781 | 0.340 | |
| Eta | 0.318 | 0.013 | 0.198 | 0.011 | 1.484 | 0.301 | 0.001 | 4.261 | 0.404 | |
@@
5.3 Comparison at Different Average Risks
@@
In this section, we conduct a comprehensive study of different truncation sampling methods at different average Risks. As discussed in Section 4.2, this allows for a fair comparison which is independent of parameter tuning. Moreover, we provide the corresponding parameters for each truncation sampling method at different average Risks, which could serve as a user reference for these methods.
@@
As can be seen in Table 1, different truncation sampling methods are compared at the average Risk of 1, 5, and 15 respectively. As discussed in Section 4.1, our defined risk and recall metrics explicitly exclude the source of risk induced by a LLM’s capacity by design, thus similar parameter values correspond to the same risk level for most sampling methods across various model types and sizes. This exactly showcases the advantage of our evaluation being tuning-independent and sustainable to the rapid update of LLMs. Among the evaluated methods, Eta-sampling Hewitt et al. (2022) is sensitive to the changes of model type and size, which might hinder its practical significance especially at a low risk level.
@@
Regarding diversity, i.e., the average recall at the same average Risk, Adaptive sampling Zhu et al. (2024) and Mirostat Basu et al. (2021) are the best and second performers, which consistently outperform the Top-k baseline by a considerable margin. Top-p mostly exhibits inferior recall comparing to the Top-k baseline, so does Eta-sampling at the average Risk of 1. As for the stability represented by standard error of Risks, Top-k sampling reaches the best scores in most cases. In comparison, Adaptive sampling and Mirostat deliver comparable standard error of risks to Top-k sampling, whereas Top-p sampling and Eta-sampling are again inferior. Considering both diversity and stability, Adaptive sampling and Mirostat are the top 2 adaptive methods to be recommended, whereas Top-p sampling shall be the last two methods to be considered.
@@
We also show in Figure 5 that larger models of the same family have higher average recall at the same risk level comparing to the smaller ones. This conforms to the fact that larger models better captures the human text distribution. Please note that our metrics does not allow a direct comparison between different model families, mainly due to their different vocabulary sizes and tokenizers, e.g., Llama-3 has a 128,256 vocabulary size, while Llama-2 has only 32, 000 vocabulary size. Moreover, our metrics also explicitly exclude the source of risk within the optimal allowed set, which is heavily dependent on a LLM’s capacity.
@@
| Avg. Risk 1 | Avg. Risk 5 | Avg. Risk 15 | |
@@ 5.4 Validation on TruthfulQA Benchmark
Although our evaluation protocol is grounded by the thorough design process with reasonable simplifications, we would like to verify its effectiveness in the real-world scenario using the TruthfulQA Benchmark Lin et al. (2021). The evaluation results using gpt2-xl are shown in Table 2. For all the methods other than greedy decoding, we run 3 times at each average risk level and report the mean and standard deviation (parenthetical value).
@@
It can be observed that greedy decoding falls far behind sampling-based decoding strategies, which conforms to the issue of likelihood-oriented decoding discussed in Section 1, as well as the findings in recent studies Cobbe et al. (2021); Wang et al. (2023); Wang and Zhou (2024); Shi et al. (2024a). All the truncation sampling methods at the low risk level achieves lower accuracy comparing to Naive sampling, due to the over-truncation of the decoding paths. At the average risk level of 5, all the truncation sampling methods slightly improve their own accuracy. Top-k sampling, Adaptive sampling and Mirostat also reach comparable or slightly higher accuracy in comparison to Naive sampling. However, further increased average risk level (means improved average recall and thus diversity) does not benefit the performance on TruthfulQA, which is plausible. Moreover, there exists a even stronger correlation between Risk SE (Standard Error of Risks) and TruthfulQA accuracy, validating the importance of stability when evaluating an adaptive decoding method. The strong correlation between TruthfulQA accuracy and our proposed average recall as well as standard error of risks at different average Risks validate the soundness and effectiveness of our evaluation method.
6 Revisiting Existing Evaluation
In this section, we revisit the recent study (Shi et al., 2024a) by comparing sampling-based decoding methods at the same average Risks. We adopt the official implementation of Shi et al. (2024a). Following their setups, we adopt Llama-2-7B on MBPP (Austin et al., 2021), HumanEval (Austin et al., 2021) and GSM8K (Cobbe et al., 2021) to evaluate coding and math problem solving performance. Mean and standard deviation for three runs are reported in Table 3, Table 4 and Table 5, respectively.
For all the three tasks, Mirostat does not perform well in general, probably because it is based on the Zipf-law of natural language and thus not suitable for code and math tasks. Notably, our greedy decoding baseline achieves significantly lower result than reported by Shi et al. (2024a) on HumanEval. Our results should be plausible, because the instruction tuned Llama-2-7B only achieves 7.9 according to Meta-Llama Github444 https://github.com/meta-llama/llama3/blob/main/MODEL_CARD.md .
While their study concludes that deterministic methods outperform sampling methods across most tasks, our evaluation reveals that sampling methods are indeed underestimated. In contrast to the conclusion in Shi et al. (2024a), all the sampling-based decoding methods could achieve better performance than greedy decoding on HumanEval in Table 4. In addition, Top-p and eta sampling also beat greedy decoding at a low average Risk on GSM8K in Table 5. This observation underscores the challenges in parameter selection for sampling-based decoding, which is effectively addressed by our method.
| Methods | Avg. Risk 1 | Avg. Risk 5 | Avg. Risk 15 |
| Top-k | 19.70 (0.50) | 21.00 (2.30) | 20.50 (0.30) |
| Top-p | 21.50 (1.30) | 21.10 (0.40) | 21.70 (0.70) |
| Mirostat | 9.50 (0.30) | 8.80 (2.00) | 8.80 (0.40) |
| Eta | 22.10 (0.70) | 19.10 (0.40) | 19.70 (0.40) |
| Greedy | 24.00 | ||
| Methods | Avg. Risk 1 | Avg. Risk 5 | Avg. Risk 15 |
| Top-k | 5.68 (2.00) | 5.08 (1.52) | 6.50 (0.76) |
| Top-p | 3.46 (1.05) | 5.89 (0.76) | 6.52 (2.43) |
| Mirostat | 3.25 (0.76) | 4.27 (1.00) | 4.68 (0.58) |
| Eta | 2.64 (2.01) | 6.91 (1.04) | 6.10 (1.32) |
| Greedy | 2.44 | ||
| Methods | Avg. Risk 1 | Avg. Risk 5 | Avg. Risk 15 |
| Top-k | 7.56 (5.39) | 11.90 (0.80) | 11.73 (0.57) |
| Top-p | 14.13 (0.47) | 8.72 (6.18) | 11.67 (0.11) |
| Mirostat | 5.46 (0.47) | 5.74 (0.64) | 3.46 (2.10) |
| Eta | 13.72 (0.46) | 8.42 (5.54) | 11.22 (0.75) |
| Greedy | 13.19 | ||
@@ 7 Conclusion
In this work, we propose an evaluation protocol to assess the trade-off between diversity and quality of truncation sampling methods for open-ended text generation. Our evaluation enjoys the merit of being independent of parameter tuning for the curated tasks. The evaluation results also serve as a user reference for different downstream tasks.
@@ 8 Limitations
In this work, we focus on the truncation sampling methods specially designed for the open-ended text generation scenario. There exist many related decoding strategies, which aim at improving different aspects of LLMs. For example, a line of decoding strategies is proposed to alleviate hallucination or improve the reasoning ability, e.g., Dola Chuang et al. (2023), Context-aware decoding Shi et al. (2024b), Contrastive decoding O’Brien and Lewis (2023) and etc. However, these methods are beyond the scope of sampling-based decoding in this study and thus not included in the discussion. Although our study is only based on text data in English for clarity, the dataset can be extended to include other languages in the future. Due to time and resource constraints, we did not include all existing sampling-based decoding methods, such as Locally Typical Sampling Meister et al. (2023) and Min-P Sampling Nguyen et al. (2024), in our comparison. However, our benchmark is publicly available, and we plan to continuously update it with evaluations of additional methods in the future.
@@ 9 Broader Impact
Our study on the capacity of sampling methods and their appropriate parameters for open-ended text generation may further promote the application of LLMs in creative industries. There exists a potential risk that our provided findings might be abused for generating harmful or fake information. However, our study itself is neutral and the mentioned risk is a general issue that LLMs face. We call for the attention on AI-Safety in the community.
@@ A.1 Complete Record of the Experiment Runs
| Run 1 at average Risks | Run 2 at average Risks | Run 3 at average Risks | average Risks | |||||||||
@@
The scores of the individual runs on TruthfulQA benchmark are recorded in Table 6, and the means and standard errors of recalls and risks at all average Risks are listed in Table 7. Note that due to a fixed amount of computation budget, we search the corresponding parameter value for each truncation sampling method till the average risk is close enough to the predefined value, thus resulting in the variations of the average risks. However, such variations are negligible given the minor differences.
Although Top-p sampling is indeed also adaptive regarding the truncation position, we show that Top-p sampling have a inherent limitation. When a larger portion of the probability mass is concentrated in the first few tokens (this often indicates smaller entropy), a fixed cumulative probability threshold will cut a longer tail off, and vice versa. However, there’s merely a weak correlation between the entropy of the LLM’s prediction and optimal truncation values, see Figure 7.
A.2 The Advantage of Probability-Independent Metrics
In this section, we explain the practical advantages of our proposed probability-independent recall and risk metrics. As can be seen in Figure 8, the empirical distribution aligns with the by gpt2-xl predicted distribution given the same prefix in general: most of the tokens which posses high likelihood in the prediction also has a high probability based on the word frequencies of our collected CP-Trie data. However, there exists two differences: